Questions 1, 2, 3, 4, 5, 6, 7, 8 are based on the following information: An epidemiological study (hypothetical) in humans in an infected country showed the following results: 1176 humans from a village were tested for dengue using an ELISA test, 168 tested positive. Out of 168 tested positives, 100 were misclassified as positive. Similarly, 408 were also misclassified as negative by the test.

Question 1. Calculate the prevalence based on the ELISA result.

a. 14.3% b. 10% c. 13.6% d. 12.8%

Question 2. How many are true positive and true negative?

a. 68 and 600 b. 68 and 1008 c. 700 and 1008 d. 100 and 400

Question 3. What is the sensitivity of the test?

a. 14.3% b. 15.6% c. 88% d. 84%

Question 4. What is the specificity of the test?

a. 85.7% b. 88.2% c. 92% d. 99%

Question 5. What is the accuracy of the test?

a. 56.8% b. 85.7% c. 64.1% d. 98.4%

Question 6. What is the prevalence based on the true status of the disease?

a. 56.8% b. 72.0% c. 69.2% d. 14.3%

Question 7. What is the positive predictive value of the test?

a. 40.5% b. 59.5% c. 42.2% d. 58.4%

Question 8. What is the negative predictive value of the test?

a. 59.5% b. 40.5% c. 56.8% d. 69.2%
Answer #1

1. Prevalence rate = number of cases detected / number of total population who were tested for dengue using ELISA test ×100.

Total number of population is 1176

Total number of detected cases is 168

So, according to the formula = 168/1176 × 100 = 14.28 ~14.3%

A ) is the correct answer.

2. 100 were misclassified as positive means in reality they were not detected as positive in the test and also they don't have the disease. So, the total purely positive is 168-100 = 68 . True positive = 68.

408 were misclassified as negative means they were wrongly put in the negative category , but actually they were not negative .

So, total positive now were 168 +408= 576.

Then ,true negative will be 1176-576= 600

A) is the correct answer.

3. Sensitivity = true positive / true positive +false negative

True positive = 68

False negative = 408

Then according to the formula =